OpenStudy (anonymous):

One type of candy sells for $2 per pound, and another type sells for $3 per pound. An order for 18 pounds of candy costs $47. How much of each type of candy was bought?

OpenStudy (kainui):

So can you think of a way to make an equation for the total price of candy and the total weight of the candy if the weight of candy 1 is X and the weight of candy 2 is Y?

OpenStudy (kainui):

It might help to list out all the information you know and try to find what you don't know. For instance you have: Total weight of candy: 18 pounds Candy 1 per weight: 2 dollars/pound etc...

OpenStudy (anonymous):

uummmmm

OpenStudy (anonymous):

no I can't

OpenStudy (kainui):

Start out small, tell me what 4 values the problem gives you. I already gave you 2.

OpenStudy (anonymous):

2/1b, 3/1b, 18 pounds, $47 dollars.

OpenStudy (kainui):

Alright, so if candy 1 is 2 dollars per pound, then that means that 1 pound of that candy is 2 dollars, right? 2 pounds of that same candy would be 4 dollars, and 3 pounds would be 6 dollars. So the total price is based off of the weight of the candy. So if we say the weight of candy 1 is X, then we can say 2x=P is the price of just the first candy. Can you come up with an equation that describes the weight of the other candy based on its weight if you use Y instead of X?

OpenStudy (anonymous):

3y=P

OpenStudy (kainui):

Boom, perfect. Now you know the total price, 47 dollars. That's the sum of both of those prices. So now are you able to make an equation for the total price and total weight?

OpenStudy (anonymous):

what do I do?

OpenStudy (kainui):

If not, that's ok, I'm just trying to help you understand because it'll be harder to do this on a test.

OpenStudy (kainui):

So 2x is the price of candy 1 and 3y is the price of candy 2 and you're given the total price, 47 dollars. 2x+3y=47 now what are x and y? The weights, and remember you were given the total weight of the candy is 18. Can you make an equation for the total weight of the candy using x, y, and 18?

OpenStudy (anonymous):

i have no idea

OpenStudy (kainui):

x is the weight of one candy, y is the weight of the other candy and 18 is the total weight of both candies combined. Think about it a little bit, how can you make an equation out of those 3 things?

OpenStudy (anonymous):

I don't know

OpenStudy (kainui):

Which equation looks right to you: x-y=18 x+y=18 xy=18 x/y=18

OpenStudy (kainui):

What problem are you having here? I'm not going to help you anymore if you don't tell me what's so hard about this.

OpenStudy (anonymous):

x/y=18?

OpenStudy (kainui):

You think the weight of one candy divided by the weight of the other candy is the total weight? Why?

OpenStudy (kainui):

If I give you 10 pounds of one candy and 20 pounds of another candy, is the total weight of the candy a half pound or 30 pounds?

OpenStudy (anonymous):

30

OpenStudy (kainui):

Write down the equation you used to figure that out and post it.

OpenStudy (kainui):

Instead of 10 and 20 pounds use x and y and instead of 30 pounds use 18

OpenStudy (anonymous):

brb...dont go away

OpenStudy (kainui):

Kk I'll be here =D

OpenStudy (anonymous):

ok. 2x+3y=47

OpenStudy (anonymous):

2x+3x=18

OpenStudy (kainui):

2x+3y=47 that's the equation that tells you the total price. But you have 2 unknowns, so you need 2 different equations. So what is an equation for the total weight? Remember, the total weight is 18. What are the weights of the candies? candy 1 is x candy 2 is y both candies is 18 Make that into an equation. Remember in the example we did earlier, Candy 1 is 10 pounds Candy 2 is 20 pounds both candies is 30 pounds write out that as an equation

OpenStudy (anonymous):

I don't get it, I've never been good at math.

OpenStudy (kainui):

You said it yourself. 20 pounds of candy plus 10 pounds of candy is 30 pounds of candy right? so 20+10=30 But instead of 20 pounds, lets just say it's x pounds of candy, cause we don't know, and y pounds of candy for the other. Then both of those is 18 pounds. I know you can do it haha.

OpenStudy (anonymous):

all I can think of is that half of 18 is 9

OpenStudy (anonymous):

theres too many things that can =18

OpenStudy (anonymous):

I don't even care anymore so thanks for your time and sorry I'm too stupid to figure it out

OpenStudy (kainui):

x+y=18

OpenStudy (kainui):

now you have 2 equations to solve for your variables.

OpenStudy (kainui):

Just rest on it a little bit. Part of math is that sometimes it's frustrating, but eventually you'll get it if you step away from the problem and come back to it.

OpenStudy (anonymous):

Ok Im going to straight up give you the answer x=7, y=11.

OpenStudy (anonymous):

meaning there are 7 pounds of 2 candy and 11 pounds of 3$ candy.

OpenStudy (anonymous):

how do you know it's not 6 and 12?

OpenStudy (anonymous):

Now I'll explain how I solved it. So let x= Weight of 2 dollar candy and y= weight of 3 dollar candy. As the question states the total pounds of candy is x+y=18. Now thats one equation done. Lets make two equations so we can solve them both simultaneously. The price of his purchase was 47$. So the "price equation" would be 2x+3y=47. (As 2$ per pound therefore if we had 4 pounds it would cost 8 and likewise for the three dollar candy, 3*4 would = 12 so basically weight multiplied by cost will give you total cost.). Now lets solve these two equations simultaneously, 2x+3y=47 and x+y=18. Lets try getting rid of a variable by multiplying the bottom equation by 2. so now the bottom equation is 2x+2y=36 (multiplying everything in the bottom by 2). Great so now we can cancel out the x variable. as 2x-2x=0, 3y-2y=1y. So 1y=47-36. y=11 and if we plug y=11 into one of the equations lets do the easy one x+y=18 we will learn that x=7. So there we go. Problem solved.

OpenStudy (anonymous):

and its not 6 and 12 because 6*2=12 and 12*3=36. 36+12=48 but in the question it says total cost was 47 and NOT 48.

OpenStudy (anonymous):

If you need an elaboration/clarification just ask.

OpenStudy (anonymous):

Thank you. :)